OpenStudy (pawanyadav):

Sec x/1+sin^2x Integrate from 0 to π/4 Pls help in solving this

hartnn (hartnn):

so, we can try the substitution, \( \large u = \sin x\) and before that, we can do some manipulation, \(\Large \dfrac{\sec x}{(1+\sin^2 x)} = \dfrac{1}{\cos x(1+\sin^2 x)}\) \(\Large \dfrac{\cos x}{\cos^2 x(1+\sin^2 x)} = \dfrac{\cos x}{(1-\sin^2 cx)(1+\sin^2 x)}\)

hartnn (hartnn):

and now we can happily substitute u = sin x

hartnn (hartnn):

let me know if you need hint in solving \(\Large \dfrac{1}{(1-u^2)(1+u^2 )}\)

OpenStudy (pawanyadav):

Am not getting the first step where Cosx cos^2 x came from

hartnn (hartnn):

sec x = 1/cos x now i wanted cos^2 x in the denominator so I multiplied numerator and denominator by cos x

OpenStudy (pawanyadav):

Thanks. Misunderstanding

imqwerty (imqwerty):

np B)

hartnn (hartnn):

you could solve it entirely? :)

OpenStudy (pawanyadav):

Am not getting the first step where Cosx cos^2 x came from

hartnn (hartnn):

huh?